need to solve and check the equation 15 / x + 9x-7 / x+2 = 9 This is 15 over x added to 9x-7 over x+2 equals 9 I know you need to find the LCD but I keep coming up with different answers Please help
Answer by MORUPHOSUOLALE(24)

Its just a simple equation if you know each simple steps.
15/X + 9X-7/x+2=9
The best and easy way is to cross multtiply.
The equation will become- 15(x+2)+9x-7(x)=(x)(x+2}(9)
15X+30+9x^2-7x=(x^2+2X}(9)
15x+30+9X^2-7x=9x^2+18x
collecting like terms- 9x^2-9x^2+15x-7x-18x=-30
15x-25x=-30
-10x=-30
x=-30/-10
x=3
TO CHECK IF WE GOT THE RIGHT ANSWER
plug in (3) the value of x into the equation and the left hand side of the equation must equal to the rght hand side
lets see-
15/x + 9x-7/x+2=9
15/3+9(3)-7/3+2=9
5 + 27-7/5 = 9
5+20/5 = 9
5+4 = 9
9 = 9
moruph osuolale,
central piedmont college,
charlotte,nc